OpenStudy (anonymous):

Need someone to check my work: Find the dimensions for the rectangle of largest area which has its base on the x-axis and its top two vertices on the graph of the following function 4/(5+x^2) I set A=2xy so 2x(4/(5+x^2)) Then I derived everything and found the zero to be sqrt(5) and the answer to be: (sqrt(5),2/5)

OpenStudy (dape):

Everything is fine, except that you used \(2x\) for the base of the rectangle, so the dimensions for \(x\) should have a factor 2. So the dimensions for the rectangle should be \(2\sqrt{5}\times\frac{2}{5}\).

OpenStudy (dape):

the x is not times, in your notation it's \((2\sqrt{5},2/5)\)

OpenStudy (anonymous):

Thanks! Some minor errors on my part that i should've caught! Was thinking it said coordinates. Thanks for your help,really appreciate it

OpenStudy (dape):

No problem, glad I could be of help.